# Insights into the structure and kinematics of a Milky Way-like galaxy

### v2

@@ Appendix B The Basics of Fourier Series and Application to our model
An advantage of using numerical simulations for determining the number of spiral arms is that we have accurate positions for both the stellar and gaseous components, allowing us to apply techniques on the top-view column density distribution. This is otherwise challenging in observations of the Milky Way due to our inside view of the Galaxy. In order to have a better understanding of the spiral pattern present in our simulation, we use Fourier transformations, as studies of nearby galaxies (e.g. Fuchs1999) and numerical simulations (e.g. Bottema2003) have done in the past. We apply this technique to our model, both using gas and stars, to infer what spiral pattern emerges when using different tracers.
The Fourier Series provides a powerful method for approximating or representing a periodic function through a combination of simple harmonic functions (sine and cosine). This approach is particularly useful for decomposing complex waveforms into basic components, enabling a detailed analysis of structures such as the spiral arms of galaxies (e.g. Davis2012).
@@ Appendix C Spiral ridges
To extract the ridges of these sharpened images, we follow the technique described in DuranCamacho2024, where we extracted the ’skeletons’ or footprints of the spiral arms from all l-v maps in our range of simulations. We provide here a brief description of this technique, but we refer the reader to the original work for further details on the methodology.
@@ Appendix D Heliocentric kinematic maps
In a similar way to the observations by khanna2023measuring, we obtain the velocity residuals by contrasting the actual velocity fields observed in our simulation against an idealised rotation curve. We compute the idealised rotation curve by assuming perfect circular rotation, with no radial ($V_{R}$) or vertical ($V_{z}$) velocities. We then consider the tangential velocity ($V_{\phi,\mathrm{ideal}}$) as being the median velocity of all particles (stellar and gaseous) included in each radial bin. This idealised rotation curve can be seen in Fig. 16. We then obtain the velocity residual, $\Delta(-V_{\phi})$, by subtracting this idealised $V_{\phi,\mathrm{ideal}}$ from the actual $V_{\phi}$ observed for the particles within that segment. These residuals will identify non-circular motions and perturbations within the model. We focus our comparison on the trends seen on mid-plane of the galaxy (i.e. $0<z<0.25$ kpc) [^8] On their work, khanna2023measuring explore higher $z$ slices up to $z=0.75$ kpc. However, velocity trends were similar across the different $z$-regions, with less extreme motions as moving away from the Galactic plane. When looking at different heights within our model, results are near identical for all values of $z$..
@@
In Figure 7, we presented the heliocentric velocity residuals, $\Delta(-V_{\phi})$ and $\Delta V_{R}$, for the data from the Gaia DR3 study by khanna2023measuring (top), alongside our model’s. For that figure, we rotate our frame to match that used in the Gaia work, such that the $x$ axis is the line connecting the observer and the galactic centre, and the sun is at $(x,y)=(0,0)$, and we restricted our coverage and resolution of the stellar component to match that of khanna2023measuring for a direct comparison. Here we include an extended version of our heliocentric velocity residual maps (Figure 17), so that we cover the galactic centre (and thus comparable to the Gaia maps from drimmel2023). In Figure 17, we also overlay, as arrows, the different cross-sections studied in Section 5.1 and Appendix E.
@@
Overall, this exercise shows that our model - at the time and location as originally selected in DuranCamacho2024 - successfully reproduces the kinematic imprints observed in the latest Gaia results, but it also underscores the rapid evolution of the kinematical pattern over time, and the difficulty in associating any given kinematical signature to the underlying presence of a spiral arm.
@@ Appendix F Spiral arm evolution across time
We examine the connection between spiral arm formation and dissolution and the large-scale kinematics in Section 5.2. To do this, we selected a specific spiral arm segment in the top-down view of the surface density distribution of our model and track its evolution over time. Figure 9 shows the surface density of the model at $t=2.6$ Gyr, highlighting the zoomed-in region corresponding to the selected spiral arm segment. Figure 20 shows the time evolution of this arm section, during $20$ Myr in timesteps of $5$ Myr represented with each column. From top to bottom, the rows display surface density, volume density, fraction of dense gas, radial and tangential velocity residuals, and the velocity divergence. We then use these maps to characterise the evolution of these properties along the spiral arm ridge as a function of time in Fig. 10.
@@
In the first column of Fig. 20, we identify an arm segment, hereafter referred to as Section 1, located near $x\sim-2.5\,\mathrm{kpc}$, where the spiral ridge is relatively faint. The bottom panel of this column shows that the gas in this segment is mildly converging, as indicated by a mild negative velocity divergence ($\nabla\cdot\mathbf{V}<0$), shown in light pink. This flow becomes more convergent over time, and leads to the growth of the spiral arm ridge, which reaches visibly higher surface (and volume) densities in the later snapshots (columns 4 and 5), $\sim 15-20$ Myr later. This convergence is also clearly visible in the radial velocities (fourth row), where at later times, this segment sits clearly at the interface between outflowing (red) and inflowing (blue) gas.
On the other side of this spiral arm section, we identified another region (Section 2), located near $x\sim 0\,\mathrm{kpc}$, which breaks/dissolves over this 20 Myr period. This ridge starts as a high density ridge with highly converging motions, but as time evolves, it becomes increasingly divergent and breaks apart. The bottom panels of this figure show that adjacent to the location of this spiral arm ridge, there is a highly divergent zone that develops and grows with time, effectively “splitting” and spreading this spiral arm segment into a wider - lower density - feature (as can be seen in the second row). The location of the arm in the first timestep is pushed downwards (following the lower edge of the split), while the upper edge is pushed outwards (and will eventually merge with the other visible spiral arm in the figure). This “opening” of the spiral arms is not just characteristic of this particular segment, but a recurring feature of the spiral arms in the simulation - best seen in the galaxy-wide time-evolution video available on the FFOGG webpage.

### v3

@@ Appendix B The basics of fourier series and application to our model
An advantage of using numerical simulations for determining the number of spiral arms is that we have accurate positions for both the stellar and gaseous components, allowing us to apply techniques on the top-view column density distribution. This is otherwise challenging in observations of the Milky Way due to our inside view of the Galaxy. In order to have a better understanding of the spiral pattern present in our simulation, we use Fourier transformations, as studies of nearby galaxies (e.g. Fuchs and Möllenhoff, 1999) and numerical simulations (e.g. Bottema, 2003) have done in the past. We apply this technique to our model, both using gas and stars, to infer what spiral pattern emerges when using different tracers.
The Fourier Series provides a powerful method for approximating or representing a periodic function through a combination of simple harmonic functions (sine and cosine). This approach is particularly useful for decomposing complex waveforms into basic components, enabling a detailed analysis of structures such as the spiral arms of galaxies (e.g. Davis et al., 2012).
@@ Appendix C Spiral ridges
To extract the ridges of these sharpened images, we follow the technique described in Durán-Camacho et al. (2024), where we extracted the ’skeletons’ or footprints of the spiral arms from all l-v maps in our range of simulations. We provide here a brief description of this technique, but we refer the reader to the original work for further details on the methodology.
@@ Appendix D Heliocentric kinematic maps
In a similar way to the observations by Khanna et al. (2023), we obtain the velocity residuals by contrasting the actual velocity fields observed in our simulation against an idealised rotation curve. We compute the idealised rotation curve by assuming perfect circular rotation, with no radial ($V_{R}$) or vertical ($V_{z}$) velocities. We then consider the tangential velocity ($V_{\phi,\mathrm{ideal}}$) as being the median velocity of all particles (stellar and gaseous) included in each radial bin. This idealised rotation curve can be seen in Fig. 16. We then obtain the velocity residual, $\Delta(-V_{\phi})$, by subtracting this idealised $V_{\phi,\mathrm{ideal}}$ from the actual $V_{\phi}$ observed for the particles within that segment. These residuals will identify non-circular motions and perturbations within the model. We focus our comparison on the trends seen on mid-plane of the galaxy (i.e. $0<z<0.25$ kpc) [^9] On their work, Khanna et al. (2023) explore higher $z$ slices up to $z=0.75$ kpc. However, velocity trends were similar across the different $z$-regions, with less extreme motions as moving away from the Galactic plane. When looking at different heights within our model, results are near identical for all values of $z$..
@@
In Figure 7, we presented the heliocentric velocity residuals, $\Delta(-V_{\phi})$ and $\Delta V_{R}$, for the data from the Gaia DR3 study by Khanna et al. (2023) (top), alongside our model’s. For that figure, we rotate our frame to match that used in the Gaia work, such that the $x$ axis is the line connecting the observer and the galactic centre, and the sun is at $(x,y)=(0,0)$, and we restricted our coverage and resolution of the stellar component to match that of Khanna et al. (2023) for a direct comparison. Here we include an extended version of our heliocentric velocity residual maps (Figure 17), so that we cover the galactic centre (and thus comparable to the Gaia maps from Gaia Collaboration et al., 2023). In Figure 17, we also overlay, as arrows, the different cross-sections studied in Section 5.1 and Appendix E.
@@
Overall, this exercise shows that our model - at the time and location as originally selected in Durán-Camacho et al. (2024) - successfully reproduces the kinematic imprints observed in the latest Gaia results, but it also underscores the rapid evolution of the kinematical pattern over time, and the difficulty in associating any given kinematical signature to the underlying presence of a spiral arm.
@@ Appendix F Spiral arm evolution across time
We examine the connection between spiral arm formation and dissolution and the large-scale kinematics in Sect.5.2 by following a representative gaseous arm segment over time.The segment is selected from the top-down surface-density map at $t=2.6$ Gyr (Fig.9, inset box). Its subsequent evolution is shown in Fig.20 over a 20 Myr interval, sampled every 5 Myr (each column). From top to bottom, Fig.20 shows the surface density, volume density, fraction of dense gas, radial and tangential velocity residuals, and the velocity divergence. We then use these maps to characterise the evolution of these properties along the spiral arm ridge as a function of time in Fig. 10.
@@
[…]